### v1

# Improving IR-based Bug Localization with Semantics-Driven Query Reduction

Abstract
Despite decades of research, software bug localization remains challenging due to heterogeneous content and inherent ambiguities in bug reports. Existing methods such as Information Retrieval (IR)-based approaches often attempt to match source documents to bug reports, overlooking the context and semantics of the source code. On the other hand, Large Language Models (LLM) (e.g., Transformer models) show promising results in understanding both texts and code. However, they have not been yet adapted well to localize software bugs against bug reports. They could be also data or resource-intensive. To bridge this gap, we propose, IQLoc, a novel bug localization approach that capitalizes on the strengths of both IR and LLM-based approaches. In particular, we leverage the program semantics understanding of transformer-based models to reason about the suspiciousness of code and reformulate queries during bug localization using Information Retrieval. To evaluate IQLoc, we refine the Bench4BL benchmark dataset and extend it by incorporating $\approx$30% more recent bug reports, resulting in a benchmark containing $\approx$7.5K bug reports. We evaluated IQLoc using three performance metrics and compare it against four baseline techniques. Experimental results demonstrate its superiority, achieving up to 58.52% and 60.59% in MAP, 61.49% and 64.58% in MRR, and 69.88% and 100.90% in HIT@K for the test bug reports with random and time-wise splits, respectively. Moreover, IQLoc improves MAP by 91.67% for bug reports with stack traces, 72.73% for those that include code elements, and 65.38% for those containing only descriptions in natural language. By integrating program semantic understanding into Information Retrieval, IQLoc mitigates several longstanding challenges of traditional IR-based approaches in bug localization.
keywords:
Bug Localization , Hybrid , Information Retrieval , LLM , Program Semantics
journal: Journal of Systems and Software
[…]
organization=Dalhousie University, city=Halifax, state=NS, country=Canada
{highlights}
We propose a novel bug localization technique that leverages the program semantics understanding of large language models (LLMs) to identify buggy source code and reformulate search queries.
A novel technique that integrates large language models into the Information Retrieval (IR)-based bug localization and leverages their contextual reasoning.
We refined and enhanced the Bench4BL dataset by adding $\approx$30% more recent bug reports (up to September 2024), bringing the total to $\approx$7.5k bug reports.
1 Introduction
Software maintenance claimed \$2.4 trillion in 2022 from the US economy, a $\approx$84% increase over the previous two years [1]. A significant factor contributing to this increase in costs is software bug, which prevents a software system from working correctly [2]. These bugs not only hinder the functionality of software but also can lead to severe system failures (e.g., AT&T Mobility outage)[3, 4, 5, 6]. Consequently, software developers dedicate 35-50% of their time tackling these bugs [7, 8]. Zou et al. [9] recently conduct a practitioner survey where they select ten different tasks related to software bugs and collect responses from 327 software practitioners (e.g., developers, project managers, testers) affiliated with leading IT companies (e.g., Google, Meta, Microsoft, Amazon). According to their survey, 82.4% of the practitioners consider bug localization as the most important or an important task. The task involves identifying the specific locations in the software code responsible for a software bug or failure.
During bug localization, developers often rely on bug reports to find the locations of problematic code. However, understanding bugs from their description in the reports and locating them in the software code is challenging and time-consuming, even for seasoned developers [10]. Bug reports contain natural language text and program artifacts such as stack traces, program elements, and code diffs [11]. The presence of these diverse contents in bug reports and the inherent ambiguities in natural language text [12] can significantly complicate the process of bug localization. To address these challenges, researchers have been actively pursuing automated solutions for bug localization over the last 50 years [13, 14, 15, 16, 17, 18, 19, 20].
Over the last few decades, many methods have been employed to automatically localize software bugs. One frequently used method is Information Retrieval (IR). Existing IR-based approaches [13, 14, 15, 16] often use pre-processed texts from bug reports as search queries and attempt to match them with source code, overlooking the context or semantics of the code. This can lead to spurious matching between bug reports and source code, given the variability of natural language text describing the software bug [21, 12]. Several existing approaches incorporate code change history, version control history, or even code authoring history to improve bug localization [22, 20]. Although additional contexts of a software bug are captured, these approaches rely primarily on statistical chances rather than a comprehensive understanding of the program semantics, limiting their effectiveness in bug localization. A recent work [23] shows that these techniques perform comparably despite incorporating additional contextual information. Furthermore, another recent work [10] suggests that the performance of existing IR-based techniques might be significantly affected by their selected queries from the bug reports.
Recently, several approaches have focused on making appropriate search queries based on bug reports to improve bug localization. Rahman and Roy [24] leverage both co-occurrences and syntactic dependencies among words from a bug report to capture meaningful search queries. In a recent work [11], the authors also show the use of static and hierarchical dependencies among program elements to construct queries. Other techniques [25, 26] make queries from a bug report through keyword expansion or noise removal. Although effective in making queries, these techniques could be limited by the vocabularies of bug reports, which are of varying quality [10]. This limitation restricts their capacity to comprehend contextual intricacies and locate defective source code documents. In contrast, deep learning-based techniques show promise in understanding the conceptual relationship between code and text [27, 28, 29]. Large Language Models (LLMs) advance this capability by leveraging transformer architectures and self-attention mechanisms to capture nuanced semantic relationships. Being trained on corpora beyond bug reports, these techniques can draw inferences from a larger knowledge base. However, they fail to leverage the benefits that the existing methods (e.g., IR-based) have to offer during bug localization. Moreover, they require extensive data and computational power, hindering their large-scale, sustainable adoption.
To design a comprehensive solution capitalizing on the strengths of both approaches above, we propose IQLoc, a novel hybrid approach combining Information Retrieval (IR) with transformer-based models to localize software bugs. Textual relevance alone may broaden a search space, potentially delivering unrelated results. Our approach narrows down the search space by leveraging program semantics understanding of the pre-trained language models during bug localization. First, IQLoc retrieves the top K source documents textually relevant to a bug report (a.k.a., query) using an IR-based approach (e.g., the BM25 algorithm [30]). Second, it trains a transformer-based, cross-encoder model to assess their relevance (based on their program semantics) to the bug report and narrows down the search space. Finally, IQLoc reduces the search query leveraging the above documents, reranks the documents using the query, and returns the suspicious source code documents.
We selected an existing benchmark dataset – Bench4BL [23] – for our experiments and extended it with recent bug reports (i.e., submitted until September 2024). To refine the dataset, we selected the bug reports that include version information and have corresponding relevant documents in their respective GitHub repositories. We further expanded it with $\approx$30% more recent bug reports, resulting in a final set of 7,483 bug reports. To evaluate our proposed technique, we employed three appropriate and widely used metrics: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K. We compare our technique with four appropriate baselines from literature – BLUiR [19], Blizzard [11], DNNLoc [17] and RLocator [31]. Across various measures, IQLoc consistently outperformed these techniques, with improvements of up to 58.52% and 60.59% in MAP, 61.49% and 64.58% in MRR, and 69.88% and 100.90% in HIT@K for the test bug reports with random and time-wise splits, respectively. Additionally, IQLoc improves bug localization over baselines in terms of MAP and MRR, achieving 91.67% and 86.49% for bug reports with stack traces, 72.73% and 66.67% for those containing program artifacts, and 65.38% and 64.29% for natural language-only bug reports.
Thus, this research makes the following contributions:
  • A novel hybrid bug localization technique, IQLoc, that capitalizes on the strengths of both traditional (e.g., query reformulation, IR) and deep learning-based approaches (e.g., transformer models) and leverages textual relevance, program semantics relevance, and language models in bug localization.
  • A refined and extended Bench4BL dataset incorporating recent bug reports, resulting in $\approx$7.5K bug reports.
  • A comprehensive evaluation of IQLoc using three popular metrics, two different splits of the dataset, and comparison with four baseline techniques [19, 11, 17, 31] from the literature.
  • A replication bundle [^1] https://github.com/asifsamir/IQLoc comprising a prototype, a carefully curated experimental dataset, configuration details, and trained and pre-trained models for third-party use and replication.
Figure 1: Buggy Code and Method Context
Table 1: An Example of Bug Report and Search Queries
Bug ID # 1416 (Spring Webflow) Rank
Bug Title Turn off snapshot creation when max-snapshots=0 72
Bug Description Creating snapshots involves serializing and compressing the flow execution object, which can lead to issues with some PersistenceContext providers. Having the ability to turn off snapshot creation will provide an option when such issues occur at the cost of losing back button support. 57
Baseline query Bug Title + Bug Description 53
IQLoc persistencecontext snapshot involves creation losing support flow object turn execution ability providers serializing 1
2 Motivational Example
In this section, we demonstrate the effectiveness of our approach –IQLoc– in bug localization using a motivating example. We present an example where our approach outperforms the baseline methods. Table 1 shows an example bug report from Spring Workflow (SFW) and Fig. 1 shows the associated buggy code. When the pre-processed version of the bug report is executed as a query by an IR-based search engine (e.g., Elasticsearch), the first buggy document is found at the 53rd position, which is not an ideal result. As shown in Table 1, the pre-processed version of the title or description fields also does not make a good query.
The example bug report discusses an issue involving snapshot creation, which could affect several classes including PersistentContext. Interestingly, the root cause of the bug involves the serialization/compression of the FlowExecution object. However, traditional techniques such as BLUiR and Blizzard consistently select PersistenceContext as a query keyword since it is found in the bug report. This leads to poor retrieval performance and they retrieve their first buggy documents at the 13th and 18th positions, respectively. In contrast, the seminal work on deep learning-based bug localization –DNNLoc– considers multiple features (e.g., textual relevance, change history, collaborative filtering), but performs poorly by retrieving the buggy document at the 28th position.
On the other hand, IQLoc attempts to understand the context of a reported bug from the source code. In particular, its cross-encoder module learns to connect bug reports to their corresponding buggy code through transformer-based, contextual learning. The cross-encoder module of our technique was able to rank the buggy code at the 11th position. Our module achieves that by capturing the context of the bug from the bug report and making a connection with the intent of the buggy code (e.g., snapshot creation from FlowExecution). More interestingly, upon reformulating the query (a.k.a., preprocessed bug report) leveraging the IR-based and cross-encoder-based components, our technique delivers the buggy source document (e.g., Fig 1) at the 1st position, which is ideal and highly promising.
3 Background
3.1 Program Semantics and Context
Program semantics refer to the meaning or behavior of a computer program, explaining how it processes inputs, performs computations, and produces outputs [32]. They provide a structured way to reason about a program’s functionality, correctness, and execution behavior [33]. However, program semantics cannot be fully understood in isolation — they require context. Variables, functions, and control flows must be interpreted by capturing their surrounding program elements. Existing literature [34] suggests that context is not a static property but rather something that emerges dynamically from the interaction and relationships among program elements. To localize software bugs, we need to analyze not only the source code but also its semantics within its context. In particular, we leverage the Large Language Models’ ability to understand program semantics in localizing software bugs.
Figure 2: Cross-Encoder’s Architecture
3.2 Cross-Encoders
The cross-encoder [35] is a transformer-based approach used for reranking documents by enabling deep interactions between the input context and the candidate. This approach processes the context and the candidate jointly within a single transformer model. The inputs are concatenated with a special separator token (e.g., [SEP]) and passed through the transformer, allowing self-attention to capture fine-grained dependencies between them (Fig. 2).
Given an input context ctxt and a candidate cand, the transformer $T$ produces a joint representation, typically extracted from the first token of the output sequence:
$y_{\text{ctxt,cand}}=h_{1}=\text{first}(T(\text{ctxt},\text{cand}))$
During encoding, the candidate label can focus on specific features within the input context, enabling the model to identify the most relevant input features for each candidate. This property makes the cross-encoder particularly effective for reranking tasks where fine-grained relevance is warranted.
Prior studies [36, 37] have demonstrated the cross-encoder’s advantages in refining retrieved results by analyzing relationships between queries and documents. We used a cross-encoder in our first study, IQLoc, to rerank source documents based on program semantic relevance between source code and bug reports.
   (a) Fine-tuning
   (b) Inference
Figure 3: Fine-tuning and Prediction of Cross-Encoder Model
4 Methodology
In this section, we present the methodology of our proposed technique –IQLoc– for software bug localization. First, we describe the fine-tuning process of the Cross-Encoder model, shown in Fig. 3, which assesses the relevance between bug reports and source documents. Then, we explain how the fine-tuned model is integrated into the overall bug localization workflow, as illustrated in Fig. 4.
4.1 Fine-tune Cross-Encoder Model
In IQLoc, we employ a cross-encoder model to determine relevance between a bug report and a code segment (Step 3, Fig. 4). First, we fine-tune the base model of CodeBERT, a transformer-based encoder model [38], using bug-fix changes and enable it to differentiate between buggy and non-buggy instances during inference (Fig. 3a). In particular, we append feed-forward network to the pre-trained CodeBERT model to achieve classification. Due to its pre-training on a large amount of natural language text and source code, the model is well-suited for our task involving bug reports and source code documents.
To fine-tune the model, we first parse each source code document from the training corpus and extract its buggy methods by leveraging the bug-fix diffs from corresponding version control history (Step 1c, Fig. 3a). We detailed this method extraction process in Section 5.1. We also establish connection between bug reports and corresponding buggy code using appropriate heuristics [23]. Then we treat bug reports (Step 1a) and buggy code pairs (Step 1e) as positive samples in our training dataset. These pairs are fed into the cross-encoder model for training with the goal of establishing positive contextual associations between buggy methods and their respective bug reports (Step 2, Fig. 3a).
To train the model to differentiate between buggy and non-buggy samples, we also add negative pairs where each pair consists of a bug report (Step 1a, Fig. 3a) and a randomly selected method body (Step 1b-dashed) to distinguish buggy from non-buggy code by leveraging program semantics. The capability stems from CodeBERT’s pre-training on a large bimodal corpus, which equips it with rich semantic representations of code structure and intent [38], as evidenced by its performance on downstream tasks (e.g., code summarization, clone detection [39, 40, 41]). Fine-tuning further adapts these representations by reinforcing links between real-world bug reports and their corresponding buggy code, enabling the model to capture subtle semantic cues associated with buggy behavior.
Figure 4: Schematic Diagram of IQLoc: (A) Indexing & Retrieval of Documents, (B) Check Relevancy, (C) Query Reformulation, (D) Bug-Localization
4.2 Corpus Indexing
As we aim to localize software bugs using an Information Retrieval (IR)-based solution, it is essential to index all the source documents of a code base (a.k.a., corpus). In our approach, we chose Elasticsearch [42] due to its efficient handling of diverse data types, seamless integration with research tools, and robustness in document indexing. We index all source code (Step 1, Fig. 4) from the 1,578 buggy versions of 42 subject-systems from the Bench4BL dataset.
4.3 Retrieval of Potentially Buggy Source Documents
We use Elasticsearch to retrieve potentially buggy source documents by executing queries selected from bug reports (Step 2, Fig. 4). Elasticsearch employs its default Standard Analyzer to preprocess queries and retrieve the top-K results (e.g., K = 100) from the corpus. During this retrieval process, we use the default scoring function of Elasticsearch, Okapi BM25 [30], to assess the relevance of the search results. When retrieving suspicious buggy documents, we specify the particular project and version of the bug report where the bug occurred. Unlike other bug localization techniques [19, 18, 11], which consider only the most recent version of a project and may retrieve unrelated documents, our approach ensures that we localize bugs within the correct historical context. This allows us to better mimic real-world scenarios by retrieving documents from the appropriate project and version. At the end of this step, we get a list of potential buggy source documents based on their textual relevance to the bug report.
4.4 Relevance Estimation using Cross-Encoder
After obtaining results from Elasticsearch, we employ our fine-tuned cross-encoder model to estimate their program semantic relevance to the bug report (Step 3, Fig. 4). Given that the cross-encoder model was trained to differentiate between buggy and bug-free code (Fig. 3.2), its reasoning against the candidate methods should be useful. While source code vocabulary is important, we thus also analyze code-level contexts leveraging the cross-encoder to support bug localization.
We analyze each retrieved source code document by parsing its Abstract Syntax Tree (AST) and extracting its individual methods (Step 1a-1b, 3b). Each method is then paired with the bug report (Step 1c, 3b) and passed into the fine-tuned cross-encoder model (Step 2, 3b), which generates generate a relevance score. This score, ranging from 0 to 1, estimates the likelihood of a method being buggy. This step refines the retrieved documents, enhancing the relevance between the bug report and source code (Step 4, Fig. 4).
4.5 Query Reformulation
Although the above steps narrow down our search space, we further refine the relevance estimation between bug reports and source code segments by reformulating our search queries. In the following section, we discuss how we reformulate our queries by leveraging the LLM’s reasoning capabilities as follows.
Algorithm 1 Extract Keywords
1:Document DocDoc, number of top keywords NN, trade-off parameter $\lambda$ ($0\leq\lambda\leq 1$)
2:procedure ExtractKeywords($Doc,N,\lambda$)
3:  $D\leftarrow\{\}$ $\triangleright$ Initialize an empty dictionary
4:  $T\leftarrow\textsc{Preprocess}(Doc)$ $\triangleright$ Get tokens
5:  for each $t\in T$ do
6:   $E\leftarrow\text{Embed}(t)$
7:   $D[t]\leftarrow E$ $\triangleright$ Add embedding for a token
8:  end for
9:  $B\leftarrow\text{Embed}(Doc)$ $\triangleright$ Compute document embedding
10:  $K\leftarrow\emptyset$ $\triangleright$ Initialize selected keyword set
11:  $C\leftarrow D.keys()$ $\triangleright$ Candidate tokens
12:  while $|K|<N$ and $C\neq\emptyset$ do
13:   $Scores\leftarrow\{\}$
14:   for each $t\in C\setminus K$ do
15:     $s_{d}\leftarrow\textsc{CosineSimilarity}(D[t],B)$ $\triangleright$ Similarity with document
16:     if $K\neq\emptyset$ then
17:      $s_{k}\leftarrow\max\limits_{k\in K}\textsc{CosineSimilarity}(D[t],D[k])$
18:     else
19:      $s_{k}\leftarrow 0$ $\triangleright$ No diversity penalty if no keywords selected
20:     end if
21:     $MMR\leftarrow\lambda\cdot s_{d}-(1-\lambda)\cdot s_{k}$ $\triangleright$ Compute MMR score
22:     $Scores[t]\leftarrow MMR$
23:   end for
24:   $t^{*}\leftarrow\arg\max\limits_{t\in C\setminus K}Scores[t]$ $\triangleright$ Select token with highest MMR score
25:   $K\leftarrow K\cup\{t^{*}\}$
26:   $C\leftarrow C\setminus\{t^{*}\}$ $\triangleright$ Remove selected token from candidates
27:  end while
28:  return $K$ $\triangleright$ Return top-N keywords
29:end procedure
4.5.1 Pre-training of Large Language Models for Software Bugs
We extract salient keywords from a bug report and relevant code segments (Steps 5a, 5b, Fig. 4), using a Transformer-based model. By leveraging the self-attention mechanism [43], Transformer models can focus on the most relevant parts of the text, understand the context, and represent the text comprehensively using high-dimensional vectors. Their representations can be further adapted to different application domains through pre-training using domain specific unstructured data [44].
To help capture salient keywords from bug reports and source code, we pre-trained a transformer model, CodeT5, using bug reports with masking. The process of collecting the pre-training dataset is explained in Section 5.1. We chose the CodeT5 model for embedding generation since it has been trained on both natural language texts and source code, which is ideal for bug reports containing various elements including text, code snippets, and program elements. Masked pre-training [45] involves randomly masking a portion of the input tokens during training, encouraging the model to learn contextual representations by predicting the masked tokens. Bug reports contain valuable information about software issues, including descriptions of bugs, code snippets, and stack traces, etc. By incorporating the masking technique to learn from bug reports, we aimed to enhance CodeT5’s general understanding of the domain-specific language, including its software engineering jargon and bug-related linguistic nuances.
During the pre-training phase, we implemented a masking mechanism where 15% of the tokens in each input sequence (i.e., bug reports) were randomly selected and substituted with a special <extra_id_XX> token. This approach aligns with the original T5 [46] model’s convention, where XX serves as a unique identifier assigned to each masked token. The identifier follows a continuous numbering system, ensuring that each masked token in the input sequence receives a distinct number in sequential order (e.g., <extra_id_0>, <extra_id_1>, <extra_id_2>, and so forth). These input sequences with the masked tokens are then fed into the model to predict the original tokens based on contextual cues. By training the model to reconstruct the original tokens from the masked input, it acquires the ability to capture the general contextual understanding of bugs from bug reports.
Keywords from Bug Reports:
Keyword extraction involves selecting a subset of words or phrases that capture the main theme of a text document, which can support various subsequent tasks (e.g., Information Retrieval [47]). For extracting keywords from bug reports, we employ EmbedRank [48] due to its simplicity and compatibility with Transformer models using the KeyBERT [49] library. The Algorithm 1 outlines the keyword extraction process. We begin by pre-processing a bug report using standard text pre-processing techniques, such as removing stop words and punctuation marks. Next, we employ our pre-trained CodeT5 model to embed each token of the bug report. Similarly, we apply the technique to embed the entire bug report. Then we calculate cosine similarity [50] to measure the semantic proximity of each token to the bug report. Based on these scores, we select the top-N most similar keywords (Step 5b, Fig. 4) for the subsequent steps. To maximize diversity in the chosen keywords, we employ the Maximal Marginal Relevance (MMR) [51] algorithm and set the MMR parameter $\lambda=0.5$, following the authors’ recommendation. Since the number of keywords affects retrieval, we also determined the optimal value of N through a controlled experiment, as discussed in $RQ_{2}$.
Keywords from Code Segments:
We follow a similar approach to extract keywords from source code segments (Step 5a, Fig. 4). However, instead of processing entire source documents, we consider only the relevant code segments identified by the cross-encoder with a confidence score above a predefined threshold (e.g., 0.5). If multiple code segments within a document are deemed relevant, we concatenate them into a single code block before applying Algorithm 1 for keyword extraction. We detail the selection of this threshold through a controlled experiment in $RQ_{3}$.
4.5.2 Reformulating the Query:
After extracting keywords from the bug report and relevant code segments, we leverage them to reformulate our queries and to enhance the retrieval (Step 6, Fig. 4). We first measure the semantic similarity between the bug report keywords and the code keywords using cosine similarity, leveraging embeddings from the CodeT5 model. Our goal was to detect the top relevant documents and capture their overlapping tokens to enhance the keywords from the bug report. This reformulated query based on such an enhancement is then used in the subsequent steps.
4.6 Bug Localization
Once we have the reformulated query, we rerank the top-K results (e.g., 100) initially retrieved from the Elasticsearch (Step 7, Fig. 4). Similar to the initial retrieval, we use the BM25 algorithm to rerank with the query constructed in the previous step. As a result, it provides a more refined set of final results (Step 8 of Fig. 4) for bug localization. Our goal was to place the buggy documents at the top positions within the ranked list through the reranking. Then the developers will encounter the buggy documents earlier and spend less time analyzing the false-positives.
5 Experiment
We evaluate our approach using the Bench4BL benchmark dataset and three appropriate performance metrics. We conduct our experiments on a cluster computing system equipped with an NVIDIA GPU with 16 GB of vRAM and compare our technique, IQLoc, against four baseline techniques from the literature. Using our experiments, we attempt to answer four research questions as follows:
  • RQ1: How does IQLoc perform in bug localization in terms of evaluation metrics?
  • RQ2: (a) How does query length affect IQLoc’s performance in bug localization? (b) What is the rationale for choosing the CodeT5 model for reformulating query? (c) Does the use of pre-trained models enhance the performance of these queries?
  • RQ3: How does the cross-encoder model perform in identifying relevant buggy source code based on program semantics?
  • RQ4: (a) Can IQLoc outperform the existing baseline techniques in bug localization? (b) How does it perform in localizing different types of bug reports compared to the baseline techniques?
Table 2: Bench4BL Dataset Summary
Project Subject-systems Bug reports
Apache 13 4,503
Jbosss 8 1,505
Spring 25 3,451
Old Subject 5 558
Total 51 10,017
5.1 Dataset Construction
We use Bench4BL [23], a benchmark dataset, for our experiments. Table 2 summarizes the Bench4BL dataset. Since the dataset is relatively old, we refine and extend it with more recent bug reports. Table 3 shows our refined and expanded dataset. In total, we spent over 80 hours refining and expanding the dataset. The following sections detail our refinement process, the collection of new bug reports, and the dataset’s split for training and testing in bug localization.
5.1.1 Refinement of Bench4BL Dataset:
The Bench4BL [23] dataset contains curated bug reports from various Java-based community projects, including Apache, JBoss, and Spring. It captures 10,017 bug reports from 51 subject systems along with their buggy/fixed versions of code and other relevant metadata. Table 2 provides a summary of the benchmark dataset.
From our initial analysis of the benchmark dataset, we noticed that many projects or bug reports lacked appropriate versioning information, making them unsuitable for our experiment. We thus dropped the bug reports without any version details and ensured that the remaining reports could be traced back to both buggy and bug-free versions of the code. During this process, we encountered several challenges, including cases where either the buggy or fixed version was missing in the code repository. We discarded the bug reports with such discrepancies, ultimately retaining a dataset of 5,753 bug reports from the three large community projects (i.e., Apache, JBoss, and Spring).
Figure 5: An Example Bug Report from JIRA
5.1.2 Expanding the Bench4BL Dataset:
After refining the original Bench4BL dataset, we expanded it by adding more recent bug reports for our experiments. In particular, we collected bug reports that were submitted by September 2024. Our expansion is limited to bug reports from the same projects and subject systems as the original dataset. These systems are managed by GitHub (e.g., Spring) and JIRA (e.g., JBoss, Apache). To collect new bug reports, we employed two distinct approaches tailored to these issue-tracking systems.
To collect bug reports from JIRA-based issue tracking systems, we used the JIRA API [52]. We also filtered issues that were of type ⓐ ‘Bug,’ had a status of ⓑ‘Resolved,’ a resolution of ⓒ ‘Fixed,’ and included both ⓓ a buggy version and ⓔ a fixed version (Fig. 5). Additionally, we considered bug reports containing explicit ⓕ Git pull request links to ensure a proper mapping between the buggy and fixed versions, reducing false positives. Once we collected the bug reports from JIRA, we used the GitHub API to extract the corresponding buggy and fixed files from the linked pull requests. Since a pull request can contain multiple commits, we considered only accepted and verified commits. When multiple commits were associated with a pull request, we included all files across these commits as a part of the fix for that bug report. This approach accounts for bug fixes that evolve over time, often involving multiple developers addressing different aspects of the same bug. Additionally, it captures cases where bug reports were closed and later reopened for further fixes. By including all relevant files, we ensure a comprehensive representation of the bug-fix process. Finally, we ensured that the versions specified in JIRA issues (e.g., semantic version, 4.8.2) matched the version tags in GitHub (e.g., camel-4.8.2) using regex-based validation, as the version strings in JIRA could differ from those in GitHub.
Figure 6: GitHub Issue Selection
Table 3: Refined and Expanded Dataset
Project Subject-systems Major versions Bug reports
Apache 11 428 3,145
Jbosss 6 303 1,416
Spring 25 847 2,922
Total 42 1,578 7,483
Figure 7: Classification of Bug Reports
On the other hand, when collecting bug reports from GitHub, we considered issues that are ⓐ ’closed’ and labeled as ⓑ ’type: bug’ (Fig. 6). To determine the version in which a bug was fixed and reduce false positives, we collected bug reports that had explicit ⓒ milestones attached, indicating the version(s) in which the bug was resolved. Since GitHub issue reports do not explicitly define the buggy version, we identified the immediate previous version tag associated with the first bug-fix commit associated with the issue in the branch tree. While this may not always correspond to the exact commit or version where the bug was introduced, it allows us to determine the latest version in which the bug exists. After that, for resolving the buggy files, we applied the same approach as used for JIRA bug reports.
Once we collected the bug reports and resolved the buggy files, we dropped the bug reports related to configuration bugs (i.e., IML, XML). Following this procedure, we complemented the original Bench4BL dataset with 1,730 recent bug reports. In total we curated 7,485 bug reports from 42 subject-systems across 1,578 buggy/fixed versions.
To gain further insight, we classify bug reports in our experimental dataset based on their content, as was done by existing literature [11]:
  • BRST (Bug Reports with Stack Traces): Includes stack traces along with text or program elements. Queries from these reports are generally noisy.
  • BRPE (Bug Reports with Program Elements): Contains program elements (e.g., method invocations, package names) but no stack traces. Queries from these reports are considered rich.
  • BRNL (Bug Reports with Natural Language Only): Lacks both program elements and stack traces. Queries from these reports are generally less informative.
To classify them, we use regular expressions adapted from the work of Rahman et al. [11]. Fig. 7 shows the classification of our curated dataset.
Table 4: Train, Validation and Test-sets
Split Type Training Validation Test Total
Random Split 5,238 748 1,497 7,483
Time-wise Split 5,236 746 1,501 7,483
Refer to caption Refer to caption
(a) Random Split (b) Time-wise Split
Figure 8: Distribution of Bug Reports in Different Dataset-Splits
5.1.3 Train and Test Set:
For our experiments, we split the dataset into training, validation, and test sets using a 70:10:20 ratio, following the standard machine learning practice for data partitioning [53, 54]. In our dataset splitting, we adopted two strategies, as was done by earlier studies [55, 56]. The first approach utilized random selection with shuffling, where the dataset was randomly divided into training and test sets. To ensure an unbiased random split, we conducted this process five times, generating five randomly shuffled datasets to facilitate separate experiments. This process delivered 5,238 bug reports for training, 1,497 for testing and 748 for validation on random trials in each of the five sets. The second strategy was a time-wise split. This type of splitting divides data into training and test sets based on their chronological order, simulating real-world scenarios where models are trained on past data and evaluated on future instances. For each subject system, we sorted the bug reports by their submission dates, split them individually into train, validation and test sets, and then combined them with similar splits from other systems. This process resulted in 5,236 bug reports for training, 1,501 for testing and 746 bug reports for validation. Table 4 summarizes our curated dataset for the experiments. Also, Fig. 8 shows how classified bug reports (i.e., BRST, BRPE, BRNL) are distributed across our training, validation and test dataset. Note, for random-split, the distribution shows the average of five random-splits.
Table 5: Cross-Encoder Dataset (Random Split)
Dataset Type Bug Reports $q,d^{+}$ Pairs (Avg.) $q,d^{-}$ Pairs (Avg.) Total (Avg.)
Training 5,238 12,694 50,776 63,470
Test 1,497 3,715 14,860 18,575
Validation 748 1,829 7,316 9,145
Table 6: Cross-Encoder Dataset (Time-wise Split)
Dataset Type Bug Reports $q,d^{+}$ Pairs $q,d^{-}$ Pairs Total
Training 5,236 12,727 50,908 63,635
Test 1,501 3,820 15,280 19,100
Validation 746 1,691 6,764 8,455
5.1.4 Expanding Training Set for Cross-Encoder:
To fine-tune our cross-encoder model, we extract both the buggy and bug-free versions of the code associated with each bug report from the GitHub repository of the respective subject system. Using a Diff tool [57], we compare between buggy and bug-free code and identify the method bodies in the buggy document that were changed to correct the bug. These altered method bodies are selected as positive instances $(q,d^{+})$ for our model training against each bug report, with a label of 1 assigned to them. This process is illustrated in Fig. 3.
In our dataset, we observed that each bug resulted in changes to a minimum of 1 and a maximum of 7 source code documents. To construct the training set for our cross-encoder, which determines the semantic relevance between a bug report and a document, we process all documents individually and generate multiple instances of positive associations. For example, if a bug report requires changes to three documents, we create three training instances, linking the altered method bodies to the bug report as positive instances for training. Our approach ensures that each training instance has a limited number of tokens, ensuring compatibility with the transformer model, which has a token limit of 512.
To generate negative samples $(q,d^{-})$, we randomly select method bodies associated with different bug reports for a given report at hand. We only consider source code from different subject systems for these negative cases. This process is reiterated four times to yield sufficient training instances for the model, and they are labeled as 0. The choice of four negative instances against each positive instance is based on a study conducted by Huang et al. [58]. They found no notable distinctions across different numbers of negative samples and also endorsed the 4:1 ratio.
For the validation and test data, we followed the same approach to generate positive and negative samples. Tables 5 and Table 6 summarize the training, validation, and test sets for the cross-encoder across two types of experiments.
5.1.5 Dataset for Pre-training Model with Bug Reports:
To pre-train an existing baseline language model (e.g., CodeT5) with domain-specific data, we collected thousands of bug reports from GitHub repositories hosting Java-based projects using the GitHub API. Our selection process involved choosing the top 100 repositories based on their star count and thus ensuring that they remained active up to the date of selection. To maintain data integrity, we targeted the issue reports labeled as ’bugs’ and collected reports submitted before April 2024. Additionally, we confirmed that none of these repositories were already included in the Bench4BL dataset to prevent any bias. After collecting the bug reports, we meticulously cleaned them to retain only their textual content using Beautiful Soup [59] and excluded repositories with fewer than five bug reports. Finally, all these steps resulted in a dataset of 70,884 bug reports from 74 repositories. Then, these bug reports were used to pre-train the CodeT5 model.
5.2 Evaluation Metrics
To evaluate IQLoc in bug localization, we use three widely used metrics- MAP, MRR, and HIT@K. These metrics have been frequently used by the relevant literature on IR-based bug localization [11, 10], and thus are highly relevant to our approach. To perform ablation study involving our cross-encoder model, we used four commonly used metrics – Accuracy, Precision, Recall, and F1 score.
Mean Average Precision (MAP)
Precision@K refers to precision for each occurrence of the buggy source document in the ranked list. Average Precision calculates the average precision@K for all buggy documents against a search query. Therefore, Mean Average Precision (MAP) is computed by averaging the AP values across all queries (Q) from a dataset.
$P_{k}=\frac{\textit{No. of Relevant Items in Top-}k}{k}$
$AP@K=\frac{1}{|D|}\sum_{k=1}^{K}P_{k}\times B_{k}$
$MAP=\frac{1}{|Q|}{\sum_{q=1}^{Q}AP@K_{q}}$
Here, $P_{k}$ calculates the precision for the $k^{th}$ element of the top $K$ items in the ranked list returned by a query. $AP@K$ computes the average precision for a list of $K$ results against a query, utilizing $B_{k}$ to determine if the $k^{th}$ document is buggy or not. $B_{k}$ outputs 1 for a match with the ground truth and 0 otherwise. $D$ represents the set of relevant instances that match the ground truth documents against a query. Finally, $MAP$ calculates the mean of the average precision ($AP@K$) across all the individual queries $q$ in the set $Q$.
Mean Reciprocal Rank (MRR)
Reciprocal Rank (RR) is associated to the rank of the first relevant result retrieved by a technique. It calculates the reciprocal of the rank of the first relevant source document within the ranked list returned by each query.
$RR_{q}=\frac{1}{\textit{Rank of First Relevant Item}}$
$MRR=\frac{1}{|Q|}\sum_{q=1}^{|Q|}RR_{q}$
Here, $RR_{q}$ represents the Reciprocal Rank for a specific query $q$. Thus, the Mean Reciprocal Rank ($MRR$) is calculated as the mean of the Reciprocal Ranks ($RR_{q}$) for all individual queries $q$ in the set $Q$.
HIT@K
HIT@K or Recall@Top-K [19] refers to the proportion of queries for which a technique returns at least one relevant document among the top $K$ retrieved results. The higher the HIT@K values, the better the performance of a bug localization technique.
$HIT@K=\frac{1}{|\mathcal{Q}|}\sum_{q=1}^{|Q|}\begin{cases}1,&r_{q}\in\mathcal{G}\\ 0,&\text{otherwise}\end{cases}$
Here, $Q$ is the set of all queries and $r_{q}$ is a binary indicator function that returns 1 if the query $q$ has at least one ground truth item $r_{q}\in G$ within the top-K results, and 0 otherwise.
Accuracy
Accuracy is a metric that determines the proportion of correctly predicted instances out of the total instances evaluated.
$\text{Accuracy}=\frac{TP+TN}{TP+FP+TN+FN}$
Here, $TP$ (True Positive) and $TN$ (True Negative) represent instances that are correctly predicted as positive and negative, respectively. Conversely, $FP$ (False Positive) and $FN$ (False Negative) represent instances that are incorrectly predicted as positive and negative, while in reality, they are negative and positive instances, respectively. In the context of bug localization, positive instances denote buggy documents, while negative instances denote non-buggy ones.
Precision
Precision measures the proportion of correctly predicted positive instances out of all positive predictions made by a model.
$\text{Precision}=\frac{TP}{TP+FP}$
Recall
Recall, also known as sensitivity, measures the proportion of correctly predicted positive instances out of all actual positive instances in the dataset.
$\text{Recall}=\frac{TP}{TP+FN}$
F1 Score
The F1 Score provides a single measure that balances between precision and recall. It is calculated as the harmonic mean of precision and recall. This balanced measure is particularly valuable in scenarios where false positives and false negatives have different implications, ensuring a robust assessment of a model’s performance.
$\text{F1 Score}=2\times\frac{Precision\times Recall}{Precision+Recall}$
Table 7: Performance of IQLoc
Split Type MAP MRR HIT@1 HIT@5 HIT@10
Random Split 0.493 0.520 0.423 0.647 0.721
Time-wise Split 0.520 0.553 0.466 0.669 0.735
Table 8: Impact of the Selection of Top-K Results from Elasticsearch
Split Type HIT@1 HIT@5 HIT@10 HIT@50 HIT@100 HIT@200
Random Split 0.389 0.619 0.703 0.843 0.889 0.929
Time-wise Split 0.396 0.618 0.822 0.904 0.937 0.961
5.3 Evaluting IQLoc
Answering RQ1: Performance of IQLoc
Table 7 presents the performance of IQLoc in terms of Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and accuracy at different top-K results (HIT@1, HIT@5, HIT@10). For the randomly split datasets, we run the experiments five times, and the table shows the mean performance from five separate runs. On the other hand, we run our experiment once for the time-wise split dataset.
In the random split, IQLoc achieves a MAP of 0.493, indicating that, on average, the relevant documents (a.k.a., buggy source code) rank higher than the irrelevant ones. Similarly, the MRR is 0.520, suggesting that the first relevant document is, on average, found within the top 2 positions. IQLoc achieves a HIT@1 of 0.423, indicating that 42.3% of the bug reports have their relevant documents (a.k.a., buggy source code) retrieved as the top-ranked result. It also achieves a HIT@5 of 0.647, indicating that $\approx$65% of the bug reports have returned at least one buggy document within the top 5 positions, whereas the HIT@10 measure is 0.721.
In the time-wise split, IQLoc demonstrates higher performance, achieving a MAP of 0.520 and an MRR of 0.553. Besides, a HIT@1 of 0.466 indicates a substantial performance improvement over its counterpart above. However, the improvement is more noticeable for HIT@5 and HIT@10 metrics, where nearly 67% of the bug reports that have at least one buggy document retrieved within the top 5 positions and 73.5% within the top 10 positions, respectively. These metrics suggest that IQLoc consistently performs well across different data splits. Moreover, its higher performance across all metrics in the time-wise split implies that IQLoc potentially captures temporal trends from past bug reports (i.e., time-wise split) and source code versions to identify recent bugs within the code.
In our approach, we capture the top 100 results retrieved by the Elasticsearch module for subsequent reranking (Step 2, Fig. 4). Previous studies in information retrieval have also used a subset of results to rerank [60, 61, 62]. Our decision was made after carefully analyzing different top-K results. As demonstrated in Table 8, for the random split test set, we observed a HIT@100 of 0.889 and a HIT@200 of 0.929. That is, for $\approx$89% of the bug reports, a relevant result can be found within the top 100 search results, and for $\approx$93% of the reports within the top 200 search results. This is a slight 4.5% increase by considering an additional 100 results. For the time-wise split, this difference is even smaller, only 2.6%. Besides, our reranking step relies on a Transformer-based cross-encoder model, which demands significant computing power. By considering only top 100 results, we thus strike a balance between the relevance of the results and the management of our computational resources.
Table 9: Performance of IQLoc for Different Classes of Bug Reports
(a) Time-wise Split
Model MAP MRR HIT@1 HIT@5 HIT@10
ST
Baseline 0.488 0.513 0.443 0.594 0.686
IQLoc 0.565 0.601 0.535 0.698 0.730
PE
Baseline 0.599 0.621 0.505 0.768 0.874
IQLoc 0.689 0.737 0.674 0.811 0.863
NL
Baseline 0.442 0.467 0.350 0.612 0.709
IQLoc 0.460 0.487 0.382 0.626 0.720
(b) Random Split
Model MAP MRR HIT@1 HIT@5 HIT@10
ST
Baseline 0.492 0.513 0.421 0.637 0.721
IQLoc 0.561 0.592 0.503 0.708 0.755
PE
Baseline 0.582 0.632 0.542 0.759 0.819
IQLoc 0.655 0.687 0.602 0.771 0.843
NL
Baseline 0.423 0.454 0.353 0.592 0.680
IQLoc 0.432 0.456 0.350 0.596 0.686
We also evaluate IQLoc’s performance in localizing different types of bug reports: bug reports containing Stack Trace (ST), Program Element (PE), and Natural Language (NL), as discussed in Section 5.1. In this case, we consider the Elasticsearch as a traditional baseline adapted from Apache Lucene [63]. Table 9 presents the results. In the time-wise split test set (Table 9a), IQLoc outperforms Elasticsearch by 15.77% and 17.15% in MAP and MRR, respectively for bug reports containing stack traces (ST). It also improves such localization by detecting at least one buggy document in the top-10 positions, with gains of 6.41%–20% in HIT@1, HIT@5, and HIT@10. For bug reports with program elements (PE), performance improvements range from 5.6% to 33.46% in the time-wise split, with a 1.2% drop in HIT@10. In contrast, our techniques improves in all metrics for the bug reports containing only natural language (NL), but the gains are smaller. IQLoc achieves only a 4.07% increase in MAP and 4.28% in MRR. HIT@K improvements range from 1.57% to 9.14%, which is lower than the gains for ST and PE. For the random split dataset, we observe a similar trend in localizing bug reports containing stack traces (ST) and program elements (PE). For ST, IQLoc outperforms Elasticsearch by 4.72%–14.02% across all metrics. For PE, the improvement ranges from 1.58% to 12.54%. However, similar to the time-wise split, performance gains for NL bug reports remain minimal, ranging from 0.44% to 2.13% across all metrics except HIT@1.
Refer to caption    Refer to caption
(a) Random Split    (b) Time-wise Split
Figure 9: Impact of Query Reduction on Retrieval Performance
We also demonstrate how our reduced queries improve bug localization through document reranking in IQLoc. To illustrate this, we present two analyses from two different perspectives. Figure 9 highlights the benefit of incorporating our reformulated queries during the reranking step by comparing two methods: baseline Elasticsearch and IQLoc, both evaluated using the top-10 results. In our analysis of randomly split test sets, we found that both techniques retrieved buggy source documents for 944 bug reports. However, IQLoc localizes at least one buggy document for 182 more bug reports that Elasticsearch could not. A similar pattern can be observed in the time-wise split dataset, where IQLoc localized 148 more bugs for which Elasticsearch could not succeed. It should be noted that baseline Elasticsearch also localized some unique bugs, but they were much lower in number.
These analysis above led us to investigate the types of bug reports for which IQLoc might excel or fail during localization. Figure 10 presents our findings for both time-wise and random splits. In the randomly split test dataset, IQLoc successfully localized the majority of bug reports containing stack traces (ST), accounting for 96.0% of the localized cases, compared to 71.1% for those containing program elements (PE) and 32.9% for bug reports classified as natural language (NL). A similar pattern is observed in the time-wise split, where 95.7% of bug reports with stack traces were localized, followed by 63.5% with PE and 40.8% with NL.
Figure 10: IQLoc’s Performance for Different Types of Bug Reports
RQ1 Summary: IQLoc demonstrates promising performance in localizing bugs, achieving a MAP score of up to 0.520— a 10.43% improvement over baseline Elasticsearch. This improvement is driven by our query reduction strategy, which uses Transformer-based reasoning of the buggy code to better match between a query and the code. Additionally, IQLoc excels at handling various types of bug reports, successfully localizing up to 96.0% of those containing stack traces.
Refer to caption Refer to caption
(a) MAP & MRR (b) HIT@K
Figure 11: Impact of Query Length on Bug Localization
Refer to caption Refer to caption
(a) MAP : CodeT5 vs. CodeBERT (b) MRR : CodeT5 vs. CodeBERT
Figure 12: Choice of Pre-trained Models for Query Reformulation
Answering RQ2: Impact of Query Length and Embedding Models on IQLoc’s Performance
Fig. 11 shows how the length of our search queries (N) influences the performance of IQLoc. Selection of search keywords is a crucial aspect of our technique since the keywords can narrow down the search scope and position the relevant documents at higher ranks during bug localization. However, determining the optimal number of keywords in a query without compromising bug localization performance poses a challenge.
To determine an optimal length for search queries (N), we use a fixed pre-trained model, the baseline CodeT5 [64] model, during document reranking step while varying the length of queries. From Fig. 11 we see that the HIT@1 increases from 0.32 with 5 keywords to a maximum of 0.455 with a keyword length of 43, representing a $\approx$42% increase. Such a trend is consistent across other metrics, with improvements of 27.64%, 37.61%, 34.61%, and 31.52% for MAP, MRR, HIT@5, and HIT@10, respectively. However, we observe diminishing returns in performance improvement for HIT@K as query length increases, reaching a plateau around the length of 15, where performance improvement slows down. To balance between performance and query specificity, we thus chose a maximum query length of 15 for our experiments.
Refer to caption Refer to caption
(a) MAP: Baseline vs. Pre-trained (b) MRR: Baseline vs. Pre-trained
Figure 13: Choice of Pre-trained, Domain-Specific Embedding Model for Query Reformulation
We also investigate the role of embedding models in our query reformulation step during bug localization, as the quality of reformulated queries directly impacts localization effectiveness. Fig. 12 compares the performance of CodeT5 and CodeBERT in terms of MAP and MRR to determine which model is more effective for generating reformulated queries in our proposed technique. Reformulated queries using CodeBERT achieve a MAP score of 0.306 for a keyword length of 5, which is 27.90% lower than CodeT5’s score of 0.392. Although increasing the query length reduces the performance gap, CodeBERT does not surpass CodeT5. For instance, at a keyword length of 43, CodeBERT still performs 3.71% lower than CodeT5. A similar trend is observed for MRR (Fig. 12b), where CodeBERT lags by 27.35% at a keyword length of 5. Based on these performance differences, we selected CodeT5 for embedding generation in our keyword extraction module, as it leads to more effective query reformulations and thus enhances the overall bug localization performance.
We also wanted to investigate if domain-specific embedding helps in constructing queries. Recognizing the importance of domain-specific pre-training to capture nuanced language features [44], we pre-trained the CodeT5-small [64] model on a comprehensive dataset containing bug reports (see Section 4.5.1 for details). This pre-training aimed to equip the model with a deeper understanding of software bugs from Java-based systems. Our findings indicate that the pre-trained model consistently outperforms the baseline model across all metrics, achieving superior performance even with fewer keywords. From Fig. 13, we see that the MAP improves from 0.39 to 0.42, representing a 9.68% increase over the baseline model using 5 keywords. Similarly, it increases from 0.45 to 0.48 with 15 keywords, indicating a 6.98% improvement. This trend holds true across various query lengths experimented with, ranging from 5 to 43 for all metrics. Therefore, we choose the CodeT5 model pre-trained with bug reports for our technique - IQLoc.
RQ2 Summary: Reformulated queries improved the performance of IQLoc in bug localization (e.g., HIT@1 by $\approx$42%) over baseline queries (i.e., bug report), with improvement maximizing on 15 keywords in a query. Similarly, a language model pre-trained with bug reports enhances the performance of IQLoc (e.g., 9.68% for MAP) by offering domain-specific embedding, with CodeT5 performing best among the models evaluated.
(a) CE Performance: Random Split
(b) CE Performance: Time-wise Split
Figure 14: Cross-Encoder’s Performance at Different Relevance Thresholds
Answering RQ3: Performance of the Cross-Encoder Model in Determining Buggy Source Code Based on Program Semantics
In this section, we investigate the effectiveness of the cross-encoder model (adopted by IQLoc) in identifying buggy code segments based on their program semantics. Transformer-based cross-encoder models are designed to output probabilistic confidence scores ranging from 0 to 1. We fine-tuned our model to classify code segments as either buggy (i.e., 1) or non-buggy (i.e., 0) against on a given bug report. Fig. 14 illustrates the performance of our adopted cross-encoder model with various configurations.
For the random split evaluation set (Fig. 14a), at threshold 0, the model achieves the lowest accuracy of 19.5%, with the lowest precision and F1 scores and the highest recall. This occurs because, at this threshold, the model classifies all instances as positive (a.k.a., buggy), capturing all true positives (TP) but also misclassifying all negative cases as false positives (FP). Since recall is calculated as $TP/(TP+FN)$ and $FN=0$ in this scenario, recall remains at 1 despite poor precision or accuracy. Once the threshold increases above 0 (e.g., 0.1), performance stabilizes, reaching an accuracy of 91.4%, with precision, recall, and F1 scores of 81.3%, 73.8%, and 75.82%, respectively. However, at threshold 1, the opposite effect occurs— the model classifies all instances as negative, correctly identifying all negative cases (TN) but misclassifying all positive cases (TP), leading to a recall of 0. Since our evaluation dataset consists of four times more negative cases than positive ones (discussed in Section 5.1), the accuracy at threshold 1 is approximately 80%. It is noteworthy that these cross-encoder results for the random split evaluation set are averaged over five independent runs.
In the time-wise split scenario (Fig. 14b), our cross-encoder model exhibits similar trends for thresholds 0 and 1. However, once the threshold surpasses 0.1, the model stabilizes and demonstrates slightly improved performance across all metrics. Starting with accuracy, the model shows a consistent upward trend, reaching $\approx$93%, which is a 1.9% improvement over the random split dataset scenario. Similarly, precision, recall, and F1 scores remain relatively stable, hovering around 88.7%, 74.9%, and 81.2%, respectively.
The precision of our cross-encoder models at both very high and low thresholds suggests that it makes predictions closer to 1 and 0 for positive (contextually relevant) and negative (contextually not relevant) outcomes, respectively, while maintaining accuracy. Given our focus on the correctness of positive predictions (i.e., precision) while maintaining overall correctness (i.e., accuracy), we set the threshold to 0.5 for our experiments, which achieves the best balance between precision and recall, as reflected in the F1 score.
RQ3 Summary: Our fine-tuned cross-encoder model can effectively identify buggy and non-buggy source code segments leveraging their program semantics and relevance to given a bug report. A threshold of 0.5 ensures an effective separation between the two types of source code by achieving up to 93% accuracy and 81.1% precision.
Answering RQ4: Comparison with the Baseline Techniques
In this section, we compare our proposed technique IQLoc against existing bug localization techniques in terms of various evaluation metrics. In particular, we compare IQLoc with five baseline techniques – Baseline Elasticsearch, BLUiR [19], Blizzard [11], DNNLoc [17], and RLocator [31].
To replicate the baseline Elasticsearch technique, we indexed all source documents of a subject system’s repository and capture the pre-processed bug reports as queries. Then, we execute the queries with the Elasticsearch engine [42], which retrieves the relevant source documents based on the BM25 algorithm [65] and Boolean query [66]. We use the default values for the parameters - k and b - in the BM25 algorithm.
BLUiR [19] employs a structured Information Retrieval approach where it leverages the structural components of both source code documents and bug reports for effective localization, which helps avoid spurious matching. To achieve this, BLUiR collects four code elements (i.e., class names, method names, variable names, and comments) from each source code document and two textual elements from each bug report (i.e., bug title and description). Then, it conducts eight searches to compute suspiciousness score for each code-text pair and then combines scores to calculate the overall suspiciousness score for a source document. For the experiments, we collected BLUiR’s replication package from the Bench4BL repository [23] and adapted it for version-based replication. It uses Indri [67], a TF-IDF [68]-based search engine, as its backend for experiments, which has become obsolete recently. Thus, we chose to replicate the technique using Apache Lucene [63] with a TF-IDF-based scoring algorithm. It should be noted that we adopted the formula suggested by Saha et al.[19] for calculating the TF and IDF metrics, which were then fed to the TF-IDF-based retrieval algorithm of Apache Lucene.
Blizzard [11] is an IR-based bug localization technique that leverages contextual information from bug reports for query construction. It categorizes bug reports into three types (i.e., NL-Natural Language, PE-Program Element, and ST-Stack Trace), employs three separate graph-based techniques to construct queries from them, and then retrieves buggy source documents from the corpus by executing the queries with the Apache Lucene engine [63]. For the replication, we collected and adapted the replication package hosted at GitHub [69] by the original authors.
DNNLoc [17] is the first technique for bug localization that combines Deep Learning and Information Retrieval. It uses several features – bug report-source code similarity (rVSM score [18]), class name similarity, collaborative filtering, bug report recency, and bug report frequency to train a deep learning model and then uses the model to localize the bugs. To replicate DNNLoc, we trained appropriate models by extracting features from our training sets, following the authors’ suggestions. During bug localization, we used these trained models to predict suspiciousness scores of source code documents and to rank them.
Table 10: Comparison between IQLoc and Baseline Techniques in Bug Localization
(a) Average Performance Metrics for Random Split
Technique MAP MRR HIT@1 HIT@5 HIT@10
Baseline Elasticsearch 0.457 0.486 0.389 0.618 0.703
BLUiR 0.470 0.504 0.393 0.655 0.745
Blizzard 0.480 0.510 0.410 0.648 0.735
DNNLoc 0.311 0.322 0.249 0.416 0.508
RLocator 0.478 0.511 0.419 0.652 0.726
IQLoc 0.493 0.520 0.423 0.647 0.721
(b) Performance Metrics for Time-wise Split
Technique MAP MRR HIT@1 HIT@5 HIT@10
Baseline Elasticsearh 0.471 0.496 0.396 0.618 0.714
BLUiR 0.508 0.539 0.429 0.686 0.786
BLIZZARD 0.494 0.525 0.419 0.670 0.753
DNNLoc 0.324 0.336 0.232 0.477 0.591
RLocator 0.505 0.532 0.439 0.683 0.731
IQLoc 0.520 0.553 0.466 0.669 0.735
RLocator [31] is a recent IR-based technique for bug localization that incorporates reinforcement learning [70], modeling the localization task as a Markov Decision Process (MDP) [71]. Before applying reinforcement learning (RL), it retrieves source documents from Elasticsearch based on the bug report and filters suspicious documents using an XGBoost [72] model. It then employs an RL agent based on an actor-critic framework [73] that attempts to rank relevant source documents at the top positions, with MAP or MRR as reward signals. We collected the replication package of RLocator from Zenodo [74] and adapted it for our study on version-based bug localization while maintaining the original specifications of the authors.
From Table 10a (random split dataset), we see that existing techniques such as baseline Elasticsearch achieve a MAP score of 0.457, while BLUiR, Blizzard, and RLocator achieve a comparative MAP score from 0.470 to 0.480, indicating $\approx$2.71%-4.89% improvement. On the other hand, DNNLoc achieves a much lower MAP score of 0.311. IQLoc, on the other hand, has a MAP score of 0.493, which is 2.71% to 58.52% better than the baseline scores. This shows that relevant documents rank higher than irrelevant ones compared to other baseline techniques. Similar improvements are observed for MRR and HIT@1, with increases of up to 61.49% and 69.88%, respectively, indicating promising performance. Although IQLoc exhibits a marginal decrease in HIT@5 and HIT@10 compared to BLUiR, Blizzard, and Rlocator, it demonstrates an improvement of up to 55.53% and 41.9% over the baseline Elasticsearch and DNNLoc measures.
In Table 10b (time-wise split dataset), IQLoc’s performance also surpasses those of baseline techniques. While Elasticsearch achieves a MAP score of 0.471, BLUiR, Blizzard, and RLocator achieve comparative MAP scores between 0.494 and 0.508; DNNLoc’s score is much lower at 0.324. In contrast, IQLoc achieves a MAP score of 0.520, representing improvements ranging from 2.36% to 60.49% over the baseline measures. Similar improvements are observed for MRR and HIT@1, ranging from 2.59%-64.58%, and 6.15%-100.9%, respectively. However, like the randomly split data (Table 10a), IQLoc experiences a marginal drop in HIT@5 and HIT@10 compared to BLUiR and Blizzard. Nonetheless, it demonstrates improvements of up to 40.25% and 24.37% over the baseline Elasticsearch and DNNLoc measures in these metrics.
Refer to caption Refer to caption
(a) Random Split
Refer to caption Refer to caption
(b) Time-wise Split
Figure 15: Comparison of Baseline Techniques in Localizing Different Types of Bugs
Refer to caption Refer to caption
(a) Random Split
Refer to caption Refer to caption
(b) Time-wise Split
Figure 16: Performance of Different Techniques on Different Subject Systems
We also demonstrate the effectiveness of IQLoc in localizing bugs from different types of bug reports compared to baseline techniques. Fig. 15 presents the performance of various methods in localizing bugs from the reports with Stack Trace (ST), Program Elements (PE), and Natural Language (NL) [11]. For the random split test set (Fig. 15a), Elasticsearch, BLUiR, and Blizzard show similar MAP scores (0.54–0.58) for ST bug reports, while RLocator performs slightly better at 0.62. IQLoc outperforms all, achieving 0.66 MAP, a 6.42% improvement over RLocator. In the case of bug reports with program elements (PE), most techniques score between 0.53 and 0.56, except Elasticsearch, which lags at 0.49, while IQLoc leads with 0.56 MAP. DNNLoc consistently underperforms, scoring 0.36 for ST and 0.33 for PE. For natural language-only (NL) bug reports, all techniques struggle, with IQLoc performing comparably. A similar trend is observed in MRR, where IQLoc improves by 6.15%–86.49% for ST and 1.72%–126.92% for PE. In the case of bug reports containing only natural language texts, MRR shows an improvement reaching $\approx$ 64%.
A similar pattern was observed in the time-wise split test set (Fig. 15b), where IQLoc achieves 7.81%–91.67% improvement in ST and 3.63%–72.73% in PE in terms of MAP. For MRR, IQLoc shows up to 80.48% improvement in bug reports with stack traces and 66.67% improvement in those with program elements. However, for bug reports containing only natural language, all techniques perform similarly, except for DNNLoc, which remains the lowest-performing model.
We further analyze each techniques performance across 42 subject systems and compare their MAP and MRR using box plots (Fig. 16). For the random split test set (Fig. 16a), IQLoc achieves a higher median MAP than that of all competitors except BLUiR. However, our technique outperforms others in terms of median MRR.
For the time-wise split test set (Fig. 16b), IQLoc outperforms all techniques with a higher median value and a more compact interquartile range (IQR) for both MAP and MRR, suggesting lower variability in performance. While BLUiR also shows a compact IQR, its median value is lower. DNNLoc, despite being less variable, consistently performs poorly across all subject systems, reinforcing its less effectiveness. Overall, IQLoc demonstrates superior performance in localizing bugs across different subject systems with a higher median measure and a reduced variability in performance.
Table 11: Statistical Test: IQLoc vs. Baselines
IQLoc vs p-value Effect-Size (Cliff’s $\delta$)
Elasticsearch 0.0098 Large (0.71)
BLUiR 0.0176 Large (0.58)
Blizzard 0.0209 Medium (0.49)
DNNLoc 0.0063 Very Large (0.78)
RLocator 0.0228 Medium (0.47)
IQLoc’s performance improvements over the baseline measures demonstrate its effectiveness across both dataset splits. To further validate this efficacy, we conducted Wilcoxon signed-rank test [75], a non-parametric statistical test. We use the SciPy library [76] to compute statistical significance of Precision@10 where we capture all buggy documents matching the ground truths within top-10 positions. For our significance analysis, we computed Precision@10 for each of the 1,501 samples across 42 subject systems in our test set, comparing IQLoc against multiple baseline approaches. During the test, we used the alternative=‘less’ parameter (i.e., baseline values are less than IQLoc values) to determine if IQLoc consistently achieved higher precision scores within the top-10 results compared to the baseline. In the case of the time-wise split dataset, IQLoc demonstrated p-values ranging from 0.0063 to 0.0228 when compared against baseline techniques (Table 11) for Precision@10. These p-values are below the significance threshold of 0.05, indicating statistical significance. Moreover, IQLoc exhibited effect sizes ranging from medium to very large, as measured by Cliff’s $\delta$ [77], further explaining the extent of differences. Note that we avoid measuring statistical significance for the randomly split dataset, as we perform five independent runs and average the results for our experiment. Given the evidence above, the null hypothesis is rejected, and IQLoc’s performance is found to be significantly higher than the baseline measures.
RQ4 Summary: IQLoc outperforms baseline techniques significantly, with MAP performance improvements of 58.52% for the random split dataset and 60.49% for the time-wise split dataset. This improvement stems from its ability to localize various types of bugs across different subject systems. Additionally, statistical significance tests confirm the superiority of our technique over baseline techniques with medium to large effect sizes.
6 Related Work
6.1 IR based Bug Localization
Bug localization has been a key area of research for decades, driven by the significant impacts and challenges of software bugs [9]. There are two primary categories of bug localization: spectra-based and Information Retrieval (IR)-based [13]. Spectra-based methods are known for their complexity and limited scalability [78, 79]. In our work, we primarily focus on Information Retrieval for bug localization.
Traditional IR-based bug localization methods [19, 18, 20, 80, 11] typically rely on the vector space model (VSM)[81], which analyzes the token overlap between bug reports and source code to identify buggy documents. Several studies have extended VSM by integrating additional contexts, such as bug report history [82], code change history [83], or version history [26], into the IR process. For instance, Zhou et al. introduced BugLocator [18] that employs a combined score of a modified VSM score and previous bug fix history to localize bugs. The traditional VSM-based scoring [81] method tends to show bias towards longer documents. To address this, they modified the VSM-based scoring and introduced rVSM scoring, which enhances the computation of textual relevance between bug reports and code elements. BLUiR [19] captures four types of structural components from the source code (i.e., class names, method names, variable names, and comments) and two components from bug reports (i.e., bug title and bug description). It then forms pairwise combinations of these components to perform eight separate searches using Indri [67], leveraging a modified TF-IDF approach. The final localization score is computed by summing the scores obtained from these individual searches. AmaLgam [20] integrates BLUiR and BugLocator techniques, along with version history inspired by Google by analyzing historical data from version control systems. These components undergo three independent systems to generate rankings before producing the final ranked lists. AmaLgam+ [80] takes a step further by incorporating stack trace and bug reporter history, alongside AmaLgam’s contexts. It ranks files using five components and then returns a final ranked list of source documents for bug localization. Recently, an IR-based technique, PathIdea [84], utilizes bug report logs (e.g., log snippets, stack traces) for bug localization. The authors employ a static analysis tool to construct a file-level call graph and reconstruct system execution paths from the logs. To determine the suspiciousness score for each file, they combine the VSM score, a log score that emphasizes files mentioned in the logs, and a path score that highlights files in the execution path. Some IR-based bug localization methods employ more complex mechanisms, such as Latent Dirichlet Allocation (LDA) [18] and Latent Semantic Index (LSI) [16]. Nonetheless, simpler methods have shown comparable performance while being more cost-efficient [23].
In our work, we employ a BM25-based approach (e.g., Elasticsearch [42]) to collect the candidate source documents. Then, our subsequent modules combine the scalability of textual relevance with Transformer-based code understanding grounded in program semantics to retrieve buggy documents at top-ranked positions.
6.2 Query Reformulation
Studies by Mills et al. [85] and Rahman et al. [10] suggest that poor queries from bug reports may adversely affect the performance of a VSM-based search. Consequently, numerous approaches propose to construct queries to assist developers in their tasks [86, 11, 87, 88]. These studies for query construction fall into two categories: frequency-based and graph-based keyword selection methods [10].
In frequency-based methods, researchers have utilized TF-IDF [68] and its variants to extract meaningful keywords from both bug reports and source code for use as queries. For instance, Gay et al. [89] employed relevance feedback (RF) from users to update the query. They implemented Rocchio’s expansion [90] method, enhancing query performance by adding terms from relevant documents with increased weights and suppressing or removing terms from irrelevant documents. Haiduc et al. [87] proposed a technique that employs a machine learning model trained on 28 query properties to recommend the best reformulation strategy for a given query. These strategies include query reduction, Rocchio expansion, RSV expansion, or Dice expansion, selected based on the query’s properties and performance.
Graph-based methods analyze semantic and syntactic dependencies among words to determine their importance. Rahman and Roy [11] applied the PageRank algorithm [91] on constructed graphs to suggest search keywords from various sources. Subsequently, they explored genetic algorithms [92] for near-optimal search query construction, leveraging the vocabulary in bug reports for effective query building [10]. However, their genetic algorithm-based approach is costly and relies solely on the textual relevance between bug reports and source documents.
These techniques typically use statistics and correlations to generate queries from bug reports. However, they often fail to capture the contextual relevance between bug reports and code which limits the effectiveness of query reformulations. In our approach, we address this by leveraging the broader understanding of natural language and program texts from large language models (e.g., CodeT5, CodeBERT) and identifying the most salient terms from both bug reports and code for query construction.
6.3 Deep Learning for Bug Localization
Recent advancements of deep learning in various domains (e.g., natural language processing, and machine translation) have encouraged its application to bug localization. DNNLOC, an early and influential work in the field of localization, is designed to identify potentially buggy documents by learning from multiple features (i.e., rVSM score, class name similarity, collaborative filtering, bug report recency, and bug report frequency) of bug reports and source code. However, its reliance on certain features like bug fixing recency and frequency is available for only 20-40% of the bugs [18]. So, the unavailability of these features might affect the performance of the model. A recent technique, FBL-BERT [93], incorporates the ColBERT model [94] for changeset-based bug localization. ColBERT uses a late interaction architecture, independently encoding bug reports and changesets with BERT, followed by efficient vector similarity calculations for relevance estimation. FBL-BERT enhances this by considering different levels of changeset granularity, enabling offline pre-computation of changeset embeddings, and employing a two-stage process of retrieving and reranking for bug localization. Nonetheless, the model may not be as effective in large-scale environments, where big software projects produces thousands of commits a day. Another recent bug localization technique, RLocator [31], is a first-of-its-kind reinforcement learning (RL) technique for bug localization. The technique attempts to optimize the ranking of a set of documents by employing RL agents based on an actor-critic [73] framework with entropy regularization, using a reward signal derived from ranking metrics (e.g., MAP and MRR) for localizing bugs. Other approaches, such as DeepLoator [95] proposed bug localization using Convolutional Neural Networks (CNN). DreamLoc [96] addresses the problem of bug localization by proposing a deep and wide architecture. The deep component applies attention mechanisms to capture textual relevance, while the wide component incorporates domain knowledge through a linear model using features such as bug-fixing history and code complexity. However, in addition to technical challenges, these techniques may encounter inefficiencies when dealing with large numbers of documents for bug localization.
In contrast, our hybrid approach addresses these challenges by focusing on a limited set of documents using Transformer models for scalability and performance. Leveraging the self-attention mechanism of the Transformer model, we process both source code and bug reports simultaneously and analyze their semantics, which helps detect their contextual relevance. This methodology offers a promising avenue for overcoming the limitations faced by existing bug localization techniques.
7 Threats to Validity
Threats to internal validity relate to experimental errors and biases. Re-implementation of the existing baseline techniques could pose a threat. For BLUiR, Blizzard and RLocator, we collected the replication packages from Bench4BL, Rahman et al.’s GitHub repositories and from Zenodo [97, 19, 74]. However, since the Indri [67] library has become obsolete, we replaced that with Apache Lucene [63] in the BLUiR replication. For DNNLOC, as the replication package is not available by the original authors, we had to replicate it ourselves. While we acknowledge the potential for implementation errors, we addressed this concern by adhering to the settings and parameters of the original authors through extensive testing. We also repeat our experiments on two different datasets [11, 18] and compare the performance with baselines to mitigate any bias due to random trials.
Threats to external validity relate to the generalizability of our work. Even though IQLoc is evaluated using only Java code, the underlying models can be adapted to different programming languages through appropriate fine-tuning [98].
Threats to construct validity relate to the evaluation metrics for our work. We use several metrics such as Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K, which are widely adopted in recommendation systems [11, 19, 99] and Information Retrieval [100, 101]. This confirms no or little threats to construct validity.
Finally, we adapted Bench4BL [23] dataset for our experiments, which might contain biases [102] and data quality issues (e.g., misclassified bugs, erroneously labeled buggy files). During the training of the Cross-Encoder model, we accepted method bodies containing buggy lines from Java classes as context. However, the impact of different context sizes was not well tested, which we consider as a scope of future work.
8 Conclusion and Future Work
Software bugs account for 50% of development time and cost billions of dollars annually. Despite significant research over the last few decades, bug localization remains a challenging task for developers. In this study, we introduce IQLoc, a novel hybrid approach that capitalizes on the strengths of both Information Retrieval (IR) and LLM-based program semantics understanding to support bug localization. Our approach enhances IR-based bug localization with reformulated queries, derived from the program understanding of Transformer models. By going beyond surface-level semantic relevance, IQLoc identifies buggy code more accurately. In our evaluation, we refined and extended the benchmark dataset Bench4BL and compared IQLoc against several baselines using three key metrics: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K. The results show that IQLoc consistently outperforms the baselines, with improvements of up to 60.59% in MAP, 64.58% in MRR, and 100.90% in HIT@K across two types of dataset splits. Additionally, further evaluation reveals that IQLoc achieves MAP improvements of 91.67% for bug reports with stack traces, 72.73% for those containing code elements, and 65.38% for those written solely in natural language. All these findings highlight the potential of IQLoc as a robust technique for bug localization that seamlessly integrates the textual relevance of traditional techniques with the program semantic understanding of modern Transformer-based models, setting a new benchmark for performance and reliability in addressing software bugs.
In the future, we plan to investigate how different representations of code structure (e.g., data flow, control flow [103]) and various tokenization strategies influence code understanding in bug localization using large language models (LLMs). Moreover, we aim to explore how agentic systems [104] can help address the multifaceted nature of bug reports in the context of bug localization using Information Retrieval (IR) techniques.